# Challenges in numerical integration in physics-informed neural networks modelling

### v2

@@ Abstract
In this paper, we numerically examine the challenges that emerge in numerical integration for various tasks now tackled by physics-informed neural networks. Specifically, we illustrate how inaccurately computed integrands can cause serious precision issues. A major difficulty lies in detecting these problems, since a simple large-scale view of the integrand may not reveal any potential errors, and the resulting outcomes are often mistakenly considered correct. To address this, it is critical to determine whether standard double-precision arithmetic suffices for evaluating the integrand or if higher precision is necessary. We evaluate and suggest numerical quadrature methods for selected examples and parameter values, particularly focusing on challenging scenarios.
Received 6 February 2025
Keywords: variable precision arithmetic; numerical integration; adaptive quadrature; physics-informed neural networks
@@ 1 Introduction
Physics-informed neural networks (PINNs) modelling involves among others a numerical calculation of integrals that usually depend on several model parameters. Optimization tasks consist of large number of integral evaluations with high precision and low computational time requirements. However, for some model parameters, many numerical quadrature algorithms fail to meet these requirements. We can observe an enormous increase in function evaluations, serious precision problems and a significant increase of computational time.
@@
The structure of the paper is as follows. In Section 2, we introduce the inaccurately evaluated functions of interest. Both a rather complicated integrand from the mathematical finance as well as a simple quartic polynomial examples are presented. In Section 3, we compare different numerical quadratures applied to studied integrals and suggest a usage the variable precision arithmetic in problematic cases. In the second part of the paper we study the heat equation with one spatial variable where the initial condition is our inaccurately evaluated function of interest. In Section 4 we show how inaccuracy influences the analytical solution and in Section 5 how it behaves in the numerical approximation of the solution calculated using PINNs. We conclude in Section 6.
2 Inaccurately evaluated functions
@@
For the purposes of presented manuscript we analyse two different examples of functions that in the models have to be numerically integrated for which an accurate evaluation is crucial. In Subsection 2.1 we start with a motivating example coming from mathematical finance that was for the first time introduced by Daněk and Pospíšil (2020). To avoid rather complicated formulas in order to clarify the main message of the present manuscript, in Subsection 2.2 we describe in detail how a simple polynomial of fourth order can be inaccurately evaluated very easily.
2.1 Inverse Fourier transform integrand example
@@
$\beta$ $=\exp\{\mu_{J}+\sigma_{J}^{2}/2\}-1.$
Figure 1: Function $f(k)$ for some values of $\sigma$.
@@ 2.2 Problematic quartic function example
Figure 4 demonstrates the inaccuracy in evaluating the function values for different values of the parameter $\delta$. Whereas a global view to the function $\phi(x)$ for values of delta less than ca $10000$ (top left picture) does not show any discrepancies, a detailed zoom (top right picture) indicates the differences in evaluating the function values in double (blue) and vpa (red) arithmetic. It is worth to remind that the differences between exact and vpa (converted back to double) values are within the machine precision. Increasing values of $\delta$ further increases the inaccuracies of the phi_double(). For $\delta$ equal to $225000$ (bottom right picture) or greater phi_double() returns zero for all arguments.
@@ 3 Numerical quadratures and their failures
The aim of this section is to show how numerical quadratures fail to numerically evaluate the integral of functions introduced in the previous section within a prescribed precision. Moreover, if adaptive quadratures are used, we observe an enormous increase of function evaluations and hence the increase of computational time.
@@
Failure of the adaptiveness of the general Gauss-Kronrod($n,2n+1$) quadrature on even a piece-wise constant function is described in Example 4.1 in Daněk and Pospíšil (2020) where authors suggest how to set the adaptive refinement in order to satisfy the local error estimate to be within the prescribed relative error, but even if this local error bound is zero, the value of the calculated integral does not have to be precise at all.
@@ 3.2 Integration of the problematic quartic function
Although the exact result ($16/15\approx 1.066666666666667$) of the simplified phi_exact() is known, many numerical quadratures fail to numerically evaluate the integral of phi_double() within a prescribed precision. In contrast to the previous example, when the further simplification of the integrand was not possible and the integration results were not satisfactory, here the automatic conversion of phi_double() to symbolic expression performed by vpaintegral gives exactly the phi_exact() and hence the exact integration result. Further we compare the integration results of quadratures integral, quad, quadl and trapz applied to integrands phi_double(), phi_vpa() and phi_exact(). In Table 2 we can see the results of numerical integration for different quadratures and different values of $\delta$. As we can see all quadratures applied to phi_vpa() or phi_exact() gives the exact results, what differs is the computational time, integral and quad being the fastest. When the quadratures were applied to phi_double() their results were problematic, worsening with increasing $\delta$. For higher $\delta$, the inaccuracy of the integrand evaluation is higher and hence the Fevals in the adaptive quadrature algorithm. In the last column we indicate the appearance of the above mentioned warnings. It is worth to mention that one has to be careful with low computational times that can sometimes occur for inaccurately evaluated integrand. The case $\delta=250000$ gives a completely wrong result $0.000000000000000$, because the integrand is already completely wrongly evaluated, see also Figure 4. Similarly as in the previous example, the massive increase of adaptive refinements is caused by the inaccurately evaluated integrand which leads to a huge increase of Fevals and consequently also the computational times, see for example the case $\delta=100000$ for quad and quadl.
@@
integral 1000 phi_double 1.066666662383775 0.287 12 No
phi_vpa 1.066666666666666 0.175 1 No
phi_exact 1.066666666666666 0.062 1 No
quad 1000 phi_double 1.066666641747667 0.133 14 No
phi_vpa 1.066666666666667 0.245 14 No
phi_exact 1.066666666666667 0.126 14 No
quadl 1000 phi_double 1.066666543018575 0.065 2 No
phi_vpa 1.066666666666667 0.100 2 No
phi_exact 1.066666666666667 0.061 2 No
trapz 1000 phi_double 1.066666666650391 0.058 1 No
phi_vpa 1.066666666666667 4.953 1 No
phi_exact 1.066666666666667 0.054 1 No
integral 10000 phi_double 1.066667217711548 0.146 11 Yes1
phi_vpa 1.066666666666666 0.230 1 No
phi_exact 1.066666666666666 0.073 1 No
quad 10000 phi_double 1.066721252633333 0.160 22 No
phi_vpa 1.066666666666666 0.237 14 No
phi_exact 1.066666666666667 0.125 14 No
quadl 10000 phi_double 1.066649867548471 0.704 128 No
phi_vpa 1.066666666666667 0.105 2 No
phi_exact 1.066666666666667 0.064 2 No
trapz 10000 phi_double 1.066667380000000 0.054 1 No
phi_vpa 1.066666666666667 5.057 1 No
phi_exact 1.066666666666667 0.056 1 No
integral 100000 phi_double 1.073588382586770 0.189 12 Yes1
phi_vpa 1.066666666666666 0.170 1 No
phi_exact 1.066666666666666 0.072 1 No
quad 100000 phi_double 1.075099781351111 25.521 5006 Yes2
phi_vpa 1.066666666666667 0.245 14 No
phi_exact 1.066666666666667 0.134 14 No
quadl 100000 phi_double 1.075775476723900 11.940 2024 Yes2
phi_vpa 1.066666666666667 0.112 2 No
phi_exact 1.066666666666667 0.071 2 No
trapz 100000 phi_double 1.071808512000000 0.062 1 No
phi_vpa 1.066666666666667 5.591 1 No
phi_exact 1.066666666666667 0.063 1 No
integral 250000 phi_double 0.000000000000000 0.082 1 No
phi_vpa 1.066666666666666 0.224 1 No
phi_exact 1.066666666666666 0.078 1 No
quad 250000 phi_double 0.000000000000000 0.104 4 No
phi_vpa 1.066666666666667 0.248 14 No
phi_exact 1.066666666666667 0.139 14 No
quadl 250000 phi_double 0.000000000000000 0.072 2 No
phi_vpa 1.066666666666667 0.137 2 No
phi_exact 1.066666666666667 0.075 2 No
trapz 250000 phi_double 0.000000000000000 0.061 1 No
phi_vpa 1.066666666666667 5.516 1 No
phi_exact 1.066666666666667 0.088 1 No
@@ 3.3 Calculation of Fourier coefficients
that are even, orthonormal at $[-1,1]$ and satisfy $g_{n}(-1)=g_{n}(1)=0$. Alternatively, an equivalent formula
@@
integral 1000 phi_double 0.891088555909676 0.148 11 No
phi_vpa 0.891088548280465 0.181 1 No
phi_exact 0.891088548280465 0.082 1 No
quad 1000 phi_double 0.891088497788002 0.210 26 No
phi_vpa 0.891088530616134 0.400 26 No
phi_exact 0.891088530616134 0.202 26 No
quadl 1000 phi_double 0.891088558410162 0.110 8 No
phi_vpa 0.891088548278784 0.191 8 No
phi_exact 0.891088548278784 0.107 8 No
trapz 1000 phi_double 0.891088548289315 0.061 1 No
phi_vpa 0.891088548280465 5.911 1 No
phi_exact 0.891088548280465 0.089 1 No
integral 10000 phi_double 0.891088948393962 0.170 11 Yes1
phi_vpa 0.891088548280465 0.181 1 No
phi_exact 0.891088548280465 0.080 1 No
quad 10000 phi_double 0.891114412663896 0.195 24 No
phi_vpa 0.891088530616134 0.400 26 No
phi_exact 0.891088530616134 0.202 26 No
quadl 10000 phi_double 0.891069033973833 0.109 8 No
phi_vpa 0.891088548278784 0.195 8 No
phi_exact 0.891088548278784 0.107 8 No
trapz 10000 phi_double 0.891089828004092 0.061 1 No
phi_vpa 0.891088548280465 5.551 1 No
phi_exact 0.891088548280465 0.063 1 No
integral 100000 phi_double 0.909820722614225 0.180 12 Yes1
phi_vpa 0.891088548280465 0.200 1 No
phi_exact 0.891088548280465 0.081 1 No
quad 100000 phi_double 0.924049764163183 29.331 5006 Yes2
phi_vpa 0.891088530616134 0.475 26 No
phi_exact 0.891088530616134 0.231 26 No
quadl 100000 phi_double 0.930107304312503 11.368 2018 Yes2
phi_vpa 0.891088548278784 0.193 8 No
phi_exact 0.891088548278784 0.107 8 No
trapz 100000 phi_double 0.907718939549998 0.061 1 No
phi_vpa 0.891088548280465 5.938 1 No
phi_exact 0.891088548280465 0.104 1 No
integral 250000 phi_double 0.000000000000000 0.089 1 No
phi_vpa 0.891088548280465 0.211 1 No
phi_exact 0.891088548280465 0.099 1 No
quad 250000 phi_double 0.000000000000000 0.106 4 No
phi_vpa 0.891088530616134 0.472 26 No
phi_exact 0.891088530616134 0.212 26 No
quadl 250000 phi_double 0.000000000000000 0.075 2 No
phi_vpa 0.891088548278784 0.192 8 No
phi_exact 0.891088548278784 0.105 8 No
trapz 250000 phi_double 0.000000000000000 0.062 1 No
phi_vpa 0.891088548280465 5.544 1 No
phi_exact 0.891088548280465 0.064 1 No

### v3

# Challenges in automatic differentiation and numerical integration in physics-informed neural networks modelling

@@ Abstract
In this paper, we numerically examine the precision challenges that emerge in automatic differentiation and numerical integration in various tasks now tackled by physics-informed neural networks (PINNs). Specifically, we illustrate how ill-posed problems or inaccurately computed functions can cause serious precision issues in differentiation and integration. A major difficulty lies in detecting these problems. A simple large-scale view of the function or good-looking loss functions or convergence results may not reveal any potential errors, and the resulting outcomes are often mistakenly considered correct. To address this, it is critical to determine whether standard double-precision arithmetic suffices or if higher precision is necessary. Three problematic use-cases for solving differential equations using PINNs are analysed in detail. For the case requiring numerical integration, we also evaluate several numerical quadrature methods and suggest particular numerical analysis steps to choose the most suitable method.
Revised 25 July, 2025
Keywords: variable precision arithmetic; automatic differentiation; numerical integration; adaptive quadrature; physics-informed neural networks
@@ 1 Introduction
Solving differential equations of physical phenomena using machine learning of neural networks (NNs) has emerged recently as a new field of scientific machine learning. Physics-informed neural networks (PINNs) are a type of universal function approximators that can embed the knowledge of any physical laws directly to the learning process (Raissi, Perdikaris, and Karniadakis, 2019). PINNs modelling tasks involve among others an automatic differentiation (Baydin, Pearlmutter, Radul, and Siskind, 2018) and a numerical calculation of integrals that usually depend on several model parameters. Optimization tasks consist of large number of integral evaluations with high precision and low computational time requirements. However, for some model parameters, many numerical quadrature algorithms fail to meet these requirements. We can observe an enormous increase in function evaluations, serious precision problems and a significant increase of computational time.
@@
The structure of the paper is as follows. In Section 2, we introduce the numerical precision challenges and give examples of inaccurately evaluated functions and their derivatives and integrals. Both a simple quartic polynomial example as well as a rather complicated function from the mathematical finance are presented. In Section 3, we compare different numerical quadratures applied to studied integrals and suggest a usage the variable precision arithmetic in problematic cases. In the second part of the paper we study three problematic use-cases, in particular one linear and one non-linear ordinary differential equation and one linear partial differential equation are introduced in Section 4 and solve them using PINNs in Section 5. We conclude in Section 6.
2 Inaccurately evaluated functions and their derivatives
@@
For the purposes of presented manuscript we introduce several numerical precision challenges in differentiation and integration of functions. In Subsection 2.2 we describe in detail how a simple polynomial of fourth order can be inaccurately evaluated very easily. Accurate evaluation of functions is crucial not only for numerical quadratures, that are in detail studied in Section 3 below, but also in differentiation. Problems occur not only in numerical differentiation, but in particular also in automatic differentiation (see Section 2.3). In Subsection 2.4 we present a more complex example coming from mathematical finance that was for the first time introduced by Daněk and Pospíšil (2020).
2.1 Numerical precision challenges
In the following, we list several important categories of problems that occur when evaluating functions and consequently in differentiation and integration. For each category we provide particular examples with short explanation.
Subtractive-cancellation (loss of significance):
  • $f(x)=1-\dfrac{\sin x}{x}$, $x\in[10^{-12},10^{-2}]$, first 8–10 digits cancel, only round-off left;
  • $f(x)=\dfrac{1-\cos x}{x^{2}}$, $x\in[0,10^{-6}]$, both numerator and denominator are of the same order and close to zero, only round-off left;
@@ 2.1 Numerical precision challenges
Extreme dynamic-range (overflow/underflow or huge ratio of scales):
  • $f(x)=\mathrm{e}^{100x}-\mathrm{e}^{100(x+\varepsilon)}$, $\varepsilon\approx 10^{-8}$, $x\gtrsim 0.7$, each term overflows to $\infty$; difference becomes $\mathrm{NaN}$;
  • $f(x)=\exp(-1/x^{2})$, $x\in[0,10^{-8}]$, underflows to 0, tail area completely missing;
[…]
Highly oscillatory or near-singular behaviour:
  • $f(x)=\dfrac{\sin(k\cdot x)}{x}$, $k\approx 10^{6}$, $0<|x|<10^{-8}$, round-off dominates,
  • $f(x)=\sin(1/x)$, $x\in[0,10^{-2}]$, infinite oscillations as $x\to 0$,
  • Bessel function of the first kind $f(x)=J_{0}(k\cdot x)$, $k\approx 10^{6}$, $x\in[0,1]$, root spacing approximately $\pi/k$ causes problems to numerical quadratures.
Apart from the above categories, one has to be careful also in other scenarios such as integrable algebraic or logarithmic singularities, piece-wise discontinuities (sharp corners), random-looking “spiky” data (e.g. noise-amplified derivatives in PINNs), very high dimension (already $d\gg 3$) or for example chaotic or stiff ODE-generated solutions (see also Sections 4.1 and 4.2 below).
@@ 2.2 Problematic quartic function example
Figure 3 demonstrates the inaccuracy in evaluating the function values for different values of the parameter $\delta$. Whereas a global view to the function $\phi(x)$ for values of delta less than ca 100001000010000 (top left picture) does not show any discrepancies, a detailed zoom (top right picture) indicates the differences in evaluating the function values in double (blue) and vpa (red) arithmetic. It is worth to remind that the differences between exact and vpa (converted back to double) values are within the machine precision. Increasing values of $\delta$ further increases the inaccuracies of the phi_double(). For $\delta$ equal to 225000225000225000 (bottom right picture) or greater phi_double() returns zero for all arguments.
@@
2.3 Numerical and automatic differentiation
Numerical approximations of derivatives are in general ill-conditioned and unstable especially due to the truncation error (often called error of the method, it is the error of approximation) and round-off error. In Figure 4 we can see the error in the forward, backward and central difference approximating the derivative of function $\sin(x)$ at the point $x_{0}=1$ for different stepsize $h$. Standard double arithmetic is used on the left and vpa on the right. In both pictures on the right we can see that error of the method is dominant, the slope of the lines (there is a $\log-\log$ scale) corresponds to the order of the truncation error ($O(h)$ for the forward and backward difference schemes and $O(h^{2})$ for the central difference). On the other hand, in double arithmetic we observe that with decreasing stepsize $h$ the round-off increases and becomes dominant.
Figure 4: Error in the difference approximation in double and vpa arithmetic.
Automatic differentiation (AD) is a computational technique for evaluating the derivatives of functions specified by computer programs, by systematically applying the chain rule to elementary operations and functions as the code executes. Unlike numerical differentiation, AD produces exact derivatives up to machine precision without step-size tuning. And whereas symbolic differentiation manipulates mathematical expressions to obtain closed-form derivative formulas, AD works directly on the program’s operations, neither requiring expression simplification nor generating unwieldy symbolic representations. By interleaving derivative computations with the original function evaluation, AD enables efficient gradient and Jacobian calculation for a wide class of functions. Although AD frameworks generate algebraically exact derivatives, they still execute those formulas in finite-precision floating-point double arithmetic. Therefore, the same numerical pathologies that occur in ordinary numerical codes re-appear in the forward or reverse passes of AD. It is worth to mention that back propagation used for training neural networks is a special case of reverse accumulation pass (Baydin, Pearlmutter, Radul, and Siskind, 2018).
2.4 Inverse Fourier transform integrand example
[…]
$\beta$ $=\exp\{\mu_{J}+\sigma_{J}^{2}/2\}-1.$
Figure 5: Function $f(k)$ for some values of $\sigma$.
@@ 3 Numerical quadratures and their failures
The aim of this section is to show how numerical quadratures fail to numerically evaluate the integral of functions introduced in the previous section within a prescribed precision. Moreover, if adaptive quadratures are used, we observe an enormous increase of function evaluations and hence the increase of computational time. This increase could be one of the first signals to start doing the numerical analysis properly, both for the quadrature used as well as checking the integrand.
@@
[…]
@@ 3.2 Integration of the problematic quartic function
Although the exact result (16/151.06666666666666716/15$\approx$ 1.06666666666666716 / 15 ≈ 1.066666666666667) of the simplified phi_exact() is known, many numerical quadratures fail to numerically evaluate the integral of phi_double() within a prescribed precision. In contrast to the previous example, when the further simplification of the integrand was not possible and the integration results were not satisfactory, here the automatic conversion of phi_double() to symbolic expression performed by vpaintegral gives exactly the phi_exact() and hence the exact integration result. Further we compare the integration results of quadratures integral, quad, quadl, trapz and tanh-sinh applied to integrands phi_double(), phi_vpa() and phi_exact(). In Table 2 we can see the results of numerical integration for different quadratures and different values of $\delta$. As we can see all quadratures applied to phi_vpa() or phi_exact() gives the exact results, what differs is the computational time, integral and quad being the fastest. When the quadratures were applied to phi_double() their results were problematic, worsening with increasing $\delta$. For higher $\delta$, the inaccuracy of the integrand evaluation is higher and hence the Fevals in the adaptive quadrature algorithm. In the last column we indicate the appearance of the above mentioned warnings. It is worth to mention that one has to be careful with low computational times that can sometimes occur for inaccurately evaluated integrand. The case $\delta=250000$ gives a completely wrong result 0.0000000000000000.0000000000000000.000000000000000, because the integrand is already completely wrongly evaluated, see also Figure 3. Similarly as in the previous example, the massive increase of adaptive refinements is caused by the inaccurately evaluated integrand which leads to a huge increase of Fevals and consequently also the computational times, see for example the case $\delta=100000$ for quad and quadl.
@@
integral 1000 phi_double 1.066666662383775 0.180 12 No
phi_vpa 1.066666666666666 0.069 1 No
phi_exact 1.066666666666666 0.035 1 No
quad 1000 phi_double 1.066666641747667 0.068 14 No
phi_vpa 1.066666666666667 0.113 14 No
phi_exact 1.066666666666667 0.063 14 No
quadl 1000 phi_double 1.066666543018575 0.040 2 No
phi_vpa 1.066666666666667 0.054 2 No
phi_exact 1.066666666666667 0.040 2 No
trapz 1000 phi_double 1.066666666650391 0.038 1 No
phi_vpa 1.066666666666667 2.346 1 No
phi_exact 1.066666666666667 0.032 1 No
tanh-sinh 1000 phi_double 1.066666609903423 0.034 1 No
phi_vpa 1.066666666666667 0.051 1 No
phi_exact 1.066666666666667 0.060 1 No
integral 10000 phi_double 1.066667217711548 0.084 11 Yes1
phi_vpa 1.066666666666666 0.073 1 No
phi_exact 1.066666666666666 0.046 1 No
quad 10000 phi_double 1.066721252633333 0.083 22 No
phi_vpa 1.066666666666666 0.176 14 No
phi_exact 1.066666666666667 0.073 14 No
quadl 10000 phi_double 1.066649867548471 0.350 128 No
phi_vpa 1.066666666666667 0.057 2 No
phi_exact 1.066666666666667 0.041 2 No
trapz 10000 phi_double 1.066667380000000 0.032 1 No
phi_vpa 1.066666666666667 2.544 1 No
phi_exact 1.066666666666667 0.033 1 No
tanh-sinh 10000 phi_double 1.066565320860938 0.031 1 No
phi_vpa 1.066666666666667 0.060 1 No
phi_exact 1.066666666666667 0.038 1 No
integral 100000 phi_double 1.073588382586769 0.125 12 Yes1
phi_vpa 1.066666666666667 0.141 1 No
phi_exact 1.066666666666666 0.091 1 No
quad 100000 phi_double 1.075099781351111 16.240 5006 Yes2
phi_vpa 1.066666666666667 0.117 14 No
phi_exact 1.066666666666667 0.072 14 No
quadl 100000 phi_double 1.075775476723900 4.734 2024 Yes2
phi_vpa 1.066666666666667 0.062 2 No
phi_exact 1.066666666666667 0.038 2 No
trapz 100000 phi_double 1.071808512000003 0.034 1 No
phi_vpa 1.066666666666667 2.355 1 No
phi_exact 1.066666666666667 0.062 1 No
tanh-sinh 100000 phi_double 1.062458555214656 0.037 1 No
phi_vpa 1.066666666666667 0.053 1 No
phi_exact 1.066666666666667 0.032 1 No
integral 250000 phi_double 0.000000000000000 0.046 1 No
phi_vpa 1.066666666666666 0.076 1 No
phi_exact 1.066666666666666 0.045 1 No
quad 250000 phi_double 0.000000000000000 0.044 4 No
phi_vpa 1.066666666666667 0.118 14 No
phi_exact 1.066666666666667 0.066 14 No
quadl 250000 phi_double 0.000000000000000 0.041 2 No
phi_vpa 1.066666666666667 0.059 2 No
phi_exact 1.066666666666667 0.038 2 No
trapz 250000 phi_double 0.000000000000000 0.037 1 No
phi_vpa 1.066666666666667 2.374 1 No
phi_exact 1.066666666666667 0.032 1 No
tanh-sinh 250000 phi_double 0.000000000000000 0.031 1 No
phi_vpa 1.066666666666667 0.052 1 No
phi_exact 1.066666666666667 0.035 1 No
@@ 3.3 Calculation of Fourier coefficients
that are even, orthonormal at [1,1][-1,1][ - 1 , 1 ] and satisfy $g_{n}(-1)=g_{n}(1)=0$. Alternatively, an equivalent formula
@@
integral 1000 phi_double 0.891088555909676 0.107 11 No
phi_vpa 0.891088548280465 0.076 1 No
phi_exact 0.891088548280465 0.042 1 No
quad 1000 phi_double 0.891088497788002 0.098 26 No
phi_vpa 0.891088530616134 0.182 26 No
phi_exact 0.891088530616134 0.098 26 No
quadl 1000 phi_double 0.891088558410162 0.065 8 No
phi_vpa 0.891088548278784 0.092 8 No
phi_exact 0.891088548278784 0.057 8 No
trapz 1000 phi_double 0.891088548289315 0.036 1 No
phi_vpa 0.891088548280466 2.369 1 No
phi_exact 0.891088548280466 0.033 1 No
tanh-sinh 1000 phi_double 0.891088505867673 0.032 1 No
phi_vpa 0.891088548280696 0.053 1 No
phi_exact 0.891088548280696 0.059 1 No
integral 10000 phi_double 0.891088948393962 0.087 11 Yes1
phi_vpa 0.891088548280465 0.073 1 No
phi_exact 0.891088548280465 0.044 1 No
quad 10000 phi_double 0.891114412663896 0.091 24 No
phi_vpa 0.891088530616134 0.181 26 No
phi_exact 0.891088530616134 0.097 26 No
quadl 10000 phi_double 0.891069033973833 0.056 8 No
phi_vpa 0.891088548278784 0.092 8 No
phi_exact 0.891088548278784 0.056 8 No
trapz 10000 phi_double 0.891089828004092 0.034 1 No
phi_vpa 0.891088548280466 2.385 1 No
phi_exact 0.891088548280466 0.033 1 No
tanh-sinh 10000 phi_double 0.890985796123688 0.032 1 No
phi_vpa 0.891088548280696 0.082 1 No
phi_exact 0.891088548280696 0.035 1 No
integral 100000 phi_double 0.909820722614225 0.094 12 Yes1
phi_vpa 0.891088548280465 0.071 1 No
phi_exact 0.891088548280465 0.043 1 No
quad 100000 phi_double 0.924049764163183 11.322 5006 Yes2
phi_vpa 0.891088530616134 0.179 26 No
phi_exact 0.891088530616134 0.099 26 No
quadl 100000 phi_double 0.930107304312503 4.773 2018 Yes2
phi_vpa 0.891088548278784 0.090 8 No
phi_exact 0.891088548278784 0.054 8 No
trapz 100000 phi_double 0.907718939549998 0.036 1 No
phi_vpa 0.891088548280465 2.376 1 No
phi_exact 0.891088548280466 0.036 1 No
tanh-sinh 100000 phi_double 0.890340539064744 0.032 1 No
phi_vpa 0.891088548280696 0.050 1 No
phi_exact 0.891088548280696 0.033 1 No
integral 250000 phi_double 0.000000000000000 0.046 1 No
phi_vpa 0.891088548280465 0.101 1 No
phi_exact 0.891088548280465 0.042 1 No
quad 250000 phi_double 0.000000000000000 0.043 4 No
phi_vpa 0.891088530616134 0.177 26 No
phi_exact 0.891088530616134 0.099 26 No
quadl 250000 phi_double 0.000000000000000 0.043 2 No
phi_vpa 0.891088548278784 0.094 8 No
phi_exact 0.891088548278784 0.054 8 No
trapz 250000 phi_double 0.000000000000000 0.035 1 No
phi_vpa 0.891088548280466 2.390 1 No
phi_exact 0.891088548280466 0.037 1 No
tanh-sinh 250000 phi_double 0.000000000000000 0.032 1 No
phi_vpa 0.891088548280696 0.059 1 No
phi_exact 0.891088548280696 0.032 1 No